John Roberts is 51 years old and has been asked to accept early retirement from his company. The company has offered John three alternative compensation packages to induce John to retire: (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

1. $187,000 cash payment to be paid immediately.
2. A 16-year annuity of $21,000 beginning immediately.
3. A 10-year annuity of $57,000 beginning at age 61.

Required:
Determine the present value, assuming that he is able to invest funds at a 6% rate, which alternative should John choose? (Round your final answers to nearest whole dollar amount.)
Answer #1

Value of option

1. $187000

2. Present value = 21000 + 21000*9.712 = $224952

3. Present value = 57000*7.360*0.558 = $234092

He should choose option 3.